# Towards Stable Machine Learning Model Retraining via Slowly Varying Sequences

### v4

@@ 2.2 A General Formulation
If Assumption 1 and Observation 1 hold, then, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all PO sequences $\bm{f^{*}}$ that solve Problem (1) using the expected losses $\mathcal{L}_{P}$ and $\mathcal{L}_{S}$, there exists a PO sequence $\bm{\hat{f^{*}}}$ that solves Problem (1) using the empirical losses ^$\hat{\mathcal{L}}_{P}$ and ^$\hat{\mathcal{L}}_{S}$, such that the following hold: $\mathcal{L}_{P}(\hat{f^{*}_{b}},D_{b})\leq\mathcal{L}_{P}(f^{*}_{b},D_{b})+2C(N,\mathcal{F},\nicefrac{{\delta}}{{B}}),\ \forall b\in[B]$ and $\mathcal{L}_{S}(\bm{\hat{f^{*}}})=\mathcal{L}_{S}(\bm{f^{*}})$.
2.3 A Tractable Restricted Formulation
@@
$\begin{split}\min_{\begin{subarray}{c}f_{b,j}\in\tilde{\mathcal{F}}_{b}\\ z_{b,j}\in\{0,1\}\end{subarray}}&\quad\sum_{b=1}^{B-1}\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}\sum_{k=1}^{|\tilde{\mathcal{F}}_{b+1}|}d(f_{b,j},f_{b+1,k})\cdot z_{b,j}\cdot z_{b+1,k}\\ \text{s.t.}&\quad\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}\mathcal{L}_{P}(f_{b,j},D_{b})\cdot z_{b,j}\leq(1+\alpha)\cdot\min_{f\in\tilde{\mathcal{F}}_{b}}\mathcal{L}_{P}(f,D_{b})\quad\forall b\in[B]\\ &\quad\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}z_{b,j}=1\quad\forall b\in[B].\end{split}$ (3)
The following theorem allows us to solve the restricted problem of selecting a slowly varying sequence of models (Problem (3)) in polynomial time:
Theorem 5.
@@
Proof.
(Proof of Theorem 5) Construct a directed graph $G$ as follows: Each node $v_{b,j}$ corresponds to model $f_{b,j}\in\tilde{\mathcal{F}}_{b}$ ($j$-th model for batch $b$). Include additional nodes $v_{\text{source}}$ and $v_{\text{sink}}$. Build directed edges from $v_{\text{source}}$ to $v_{1,j},\forall j,$ and from $v_{B,j}$ to $v_{\text{sink}},\forall j$, with zero weights. For $b\in[B-1]$, build directed edges from $v_{b,j}$ to $v_{b+1,k},\forall(k,j)$; assign a weight of $d(f_{b,j},f_{b+1,k})$ to each edge $(v_{b,j},v_{b+1,k})$. To enforce the accuracy tolerance constraint, for each batch $b$, include only those nodes $v_{b,j}$ corresponding to models $f_{b,j}$ that satisfy $\mathcal{L}_{P}(f_{b,j},D_{b})\leq(1+\alpha)\cdot\min_{f\in\mathcal{F}_{b}}\mathcal{L}_{P}(f,D_{b})$. Finding a sequence of models that minimizes the total stability metric while satisfying the accuracy tolerance constraint is equivalent to finding the shortest path from $v_{\text{source}}$ to $v_{\text{sink}}$ in graph $G$. Since the shortest path problem in a directed graph with non-negative weights can be solved in polynomial time (e.g., using Dijkstra’s algorithm), Problem (3) can also be solved in polynomial time. ∎
The restricted formulation (Problem (3)) is an approximation of the full formulation (Problem (1)). To assess the quality of the approximation and the impact of the size of $\mathcal{F}_{b}$ (number of candidate models per batch) on the restricted formulation, we train slowly varying sequences of linear regression models on synthetic data generated according to $y=\bm{x}^{\top}\bm{\beta}_{b}+\epsilon$, where, for each batch, the data $X\in\mathbb{R}^{n\times p}(\text{with }n=100,p=10)$ is drawn from $\bm{x}\sim\mathcal{N}(\bm{0},\Sigma),\epsilon\sim\mathcal{N}(0,\sigma)$ and the regression coefficients satisfy $|\bm{\beta}_{b}-\bm{\beta}_{0}|\leq 0.3$ for some $\bm{\beta}_{0}\in\{-1,0,1\}$. Note that, thanks to the convexity of the linear regression optimization problem, by varying the value of $\epsilon_{b}$ in Problem (1), we recover the full empirical Pareto frontier. This is visually confirmed by the solid lines in Figure 1, which reveals that the Pareto frontier is convex: with a small increase in predictive power loss, we can achieve a significant decrease in stability loss. In contrast, the restricted model does not guarantee recovery of PO solutions due to the finite sampling of models within each batch, potentially missing the PO sequence for certain accuracy tolerances. Therefore, as expected, the restricted model finds solutions that are Pareto-dominated by the ones of the full model, whereas increasing $|\mathcal{F}_{b}|$ improves the approximation. Next, we continue with a more quantitative discussion of the findings of Fig. 1.
As can be seen by the full model’s in-sample curve, we examine accuracy tolerances of $\alpha\in[0,0.01,0.03,0.1,0.3]$, where the predictive power validation loss is constrained to be within $\alpha$ of the best independent model for each batch: the corresponding gains in stability loss range between $2-12\%$. By comparing this curve, with the restricted models’ in-sample curves, we see that, for $\alpha\leq 0.03$ (which we most care about in practice), the restricted models are about $1-3\%$ worse in terms of both predictive power and stability. By comparing the in-sample with the out-of-sample curves for both the full and the restricted model, we empirically see Theorem 4’s guarantee: the in-/out-of-sample predictive power gap is around 20%, while the gaps between the full and the restricted models remain the same in-/out-of-sample.
@@
2.4 Model Distance Calculation
To compute the distance $d(f_{i},f_{j})$ between models $f_{i}$ and $f_{j}$, we consider two primary approaches: structural distances and feature importance-based distances.
Structural distances: This approach measures the distance by directly comparing the models’ structures. For linear/logistic regression models, we compute the distance as $d(f_{i},f_{j})=\|\bm{\beta}_{i}-\bm{\beta}_{j}\|_{2}^{2}$, where $\bm{\beta}_{i}$ and $\bm{\beta}_{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively. For decision tree-based models, we use the decision tree distance metric proposed by [27] that relies on optimally matching the paths in the two trees $\mathbb{T}_{i}$ and $\mathbb{T}_{j}$, denoted as $\mathcal{P}(\mathbb{T}_{i})$ and $\mathcal{P}(\mathbb{T}_{j})$. The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account the differences in feature ranges in the two paths as well as the class labels associated with them. Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem. Formally, assuming without loss of generality that $|\mathcal{P}(\mathbb{T}_{i})|\geq|\mathcal{P}(\mathbb{T}_{j})|$, and given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{i}$ is matched with path $q$ in $\left.\mathbb{T}_{j}\right)$, the distance is defined as the following linear optimization problem:
$\min_{\begin{subarray}{c}\quad x_{pq}\in[0,1]\\ \forall(p,q)\\ \in\mathcal{P}(\mathbb{T}_{i})\times\mathcal{P}(\mathbb{T}_{j})\end{subarray}}\sum_{p\in\mathcal{P}(\mathbb{T}_{i})}\sum_{q\in\mathcal{P}(\mathbb{T}_{j})}d(p,q)x_{pq}\ \text{s.t.}\ \sum_{q\in\mathcal{P}(\mathbb{T}_{j})}x_{pq}=1,\ \forall p\in\mathcal{P}(\mathbb{T}_{i}),\ \ \sum_{p\in\mathcal{P}(\mathbb{T}_{i})}x_{pq}\geq 1,\ \forall q\in\mathcal{P}(\mathbb{T}_{j}).$ (4)
Feature importance-based distances: This approach measures the distance indirectly by comparing the models’ respective feature importances. Denoting by $\bm{\phi}_{f}\in\mathbb{R}^{p}$ the $p$-dimensional vector of feature importances that model $f$ assigns to each feature, we compute the distance as $d(f_{i},f_{j})=\|\bm{\phi}_{f_{i}}-\bm{\phi}_{f_{j}}\|_{2}^{2}$. The advantage of using feature importance-based distances is its applicability to both interpretable and black-box models. For decision tree-based models, we compute feature importances using measures intrinsic to tree growth, examining how each feature impacts node purity or classification accuracy. In particular, for CART [33, 34], we use the Gini importance, which quantifies how much a feature contributes to homogenizing the target variable within the subsets created by the splits. Given a tree $f$ with a set of nodes $V$, the Gini importance for feature $j$ is computed as $[\phi_{f}]_{j}=\frac{\sum_{v\in V}I(j,v)}{\sum_{v\in V}I(v)}$, where $I(j,v)$ represents the Gini impurity reduction at node $v$ due to splitting on feature $j$ and $I(v)$ is the total Gini impurity reduction at node $v$ regardless of the feature. For XGBoost [35], we use Gain, which focuses on the improvement in model accuracy from splits made by each feature. For general black-box models, we use SHAP (SHapley Additive exPlanations) [36], which offer a cohesive methodology to measure feature importance across different model architectures by calculating the average marginal contribution of each feature to the prediction.
@@ 2.5 Adaptive Model Retraining
The intra-sequence stability loss is defined as the sum of pairwise distances of adjacent models of one sequence, i.e., $\mathcal{L}_{S}(\bm{f})=\sum_{b=1}^{B-1}d(f_{b},f_{b+1})$.
Definition 3.
@@
The inter-sequence stability loss is defined as the distance between the final models of two sequences of lengths $B$ and $B^{\prime}$ respectively, i.e., $\mathcal{L}_{S}(\bm{f},\bm{f^{\prime}})=d(f_{B},f^{\prime}_{B^{\prime}})$.
@@
For the inter-sequence stability loss, we could alternatively compare all pairs of models in $\bm{f},\bm{f^{\prime}}$. However, such an approach (i) does not directly apply to sequences of different lengths and (ii) ignores the fact that we mostly care about the final model from each sequence (which is used in production). The next result promises that, under certain conditions, intra-sequence stability implies inter-sequence stability. In particular, we consider the implications of training a sequence of models that minimizes the intra-sequence stability loss on the inter-sequence stability when a new batch of data is introduced, and the entire sequence is retrained. We work under the additional assumption that the initial model is common in both sequences, i.e., $f_{1}=f_{1}^{\prime}=f_{0}$. This is a realistic assumption as, in practice, $f_{0}$ can be the model that was previously used in production. Further, this assumption can easily be incorporated as an extra constraint in the proposed optimization formulation.
Theorem 6.
@@
Let $\bm{f},\bm{f^{\prime}}$ be two sequences of models trained using data $D_{1},\dots,D_{B}$ and $D_{1},\dots,D_{B},D_{B+1}$, respectively, and a stability loss metric $d:\mathcal{F}\times\mathcal{F}\mapsto\mathbb{R}_{\geq 0}$. Assume (i) their intra-sequence stability losses satisfy $\mathcal{L}_{S}(\bm{f})\leq\epsilon$ and $\mathcal{L}_{S}(\bm{f^{\prime}})\leq\epsilon^{\prime}$, and (ii) the initial models satisfy $f_{1}=f_{1}^{\prime}=f_{0}$. Then, $\mathcal{L}_{S}(\bm{f},\bm{f^{\prime}})\leq\epsilon+\epsilon^{\prime}$.
Proof.
(Proof of Theorem 6) The result follows by repeated application of the triangle inequality:
$\begin{split}\mathcal{L}_{S}(\bm{f},\bm{f\textquoteright})&:=d(f_{B},f\textquoteright_{B+1})\leq d(f_{B-1},f_{B})+d(f_{B-1},f\textquoteright_{B+1})\leq\dots\leq\sum_{b=1}^{B-1}d(f_{b},f_{b+1})+d(f_{1},f\textquoteright_{B+1})\\ &\leq\dots\leq\mathcal{L}_{S}(\bm{f})+\sum_{b=1}^{B}d(f\textquoteright_{b},f\textquoteright_{b+1})+d(f_{1},f\textquoteright_{1})\leq\mathcal{L}_{S}(\bm{f})+\mathcal{L}_{S}(\bm{f\textquoteright})+d(f_{0},f_{0})=\epsilon+\epsilon\textquoteright.\end{split}$

### v5

@@ 2.4 Pareto Excess Risk Bound
Under Assumption 2.1 and Remark 2.2, for any $\delta\in(0,1)$, with probability at least $1-\delta$, for all PO sequences $\boldsymbol{f^{*}}$ that solve Problem (1) using the expected losses $\mathcal{L}_{P}$ and $\mathcal{L}_{S}$, there exists a PO sequence $\boldsymbol{\hat{f^{*}}}$ that solves Problem (1) using the empirical losses ^$\hat{\mathcal{L}}_{P}$ and ^$\hat{\mathcal{L}}_{S}$, such that the following hold: $\mathcal{L}_{P}(\hat{f^{*}_{b}},D_{b})\leq\mathcal{L}_{P}(f^{*}_{b},D_{b})+2C(N,\mathcal{F},\frac{\delta}{B}),\ \forall b\in[B]$ and $\mathcal{L}_{S}(\boldsymbol{\hat{f^{*}}})=\mathcal{L}_{S}(\boldsymbol{f^{*}})$.
For example, for linear regression, a standard finite-sample result states that, with probability at least $1-\delta$, the excess risk of the empirical risk minimizer of batch $b$, $\hat{f}_{b}$, satisfies:
$\mathcal{L}_{P}(\hat{f}_{b},D_{b})-\mathcal{L}_{P}(f^{*}_{b},D_{b})\leq\frac{\sigma^{2}p}{N_{b}}+\sigma\sqrt{\frac{4p\log(1/\delta)}{N_{b}}}.$ (3)
This bound holds independently for each batch, ensuring that each $\hat{f}_{b}$ generalizes well to new data. In our setting, however, we seek to assess the Pareto optimality of the full sequence $(f_{1},\dots,f_{B})$. By applying Lemma 2.7, we obtain the following multi-objective extension of (3):
$\mathcal{L}_{P}(\hat{f^{*}_{b}},D_{b})-\mathcal{L}_{P}(f^{*}_{b},D_{b})\leq\frac{2\sigma^{2}p}{N_{b}}+2\sigma\sqrt{\frac{4p\log\frac{B}{\delta}}{N_{b}}},\ \forall b\in[B].$ (4)
This result guarantees that the entire sequence $(\hat{f}_{1},\dots,\hat{f}_{B})$ remains Pareto optimal with high probability. Ensuring generalization across the entire sequence results in a slightly larger excess risk bound due to the tighter confidence level $\frac{\delta}{B}$, growing logarithmically with the number of models $B$; at the same time, the increasing sample size $N_{b}$ guarantees that the empirical Pareto frontier still converges to the true Pareto frontier. We numerically study the Pareto excess risk bound for linear regression in Appendix B.
3 A Practical Algorithm for Model Retraining
3.1 Tractable Restricted Formulation
@@
$\begin{split}\min_{\begin{subarray}{c}f_{b,j}\in\tilde{\mathcal{F}}_{b}\\ z_{b,j}\in\{0,1\}\end{subarray}}&\quad\sum_{b=1}^{B-1}\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}\sum_{k=1}^{|\tilde{\mathcal{F}}_{b+1}|}d(f_{b,j},f_{b+1,k})\cdot z_{b,j}\cdot z_{b+1,k}\\ \text{s.t.}&\quad\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}\mathcal{L}_{P}(f_{b,j},D_{b})\cdot z_{b,j}\leq(1+\alpha)\cdot\\ &\qquad\cdot\min_{f\in\tilde{\mathcal{F}}_{b}}\mathcal{L}_{P}(f,D_{b})\quad\forall b\in[B]\\ &\quad\sum_{j=1}^{|\tilde{\mathcal{F}}_{b}|}z_{b,j}=1\quad\forall b\in[B].\end{split}$ (5)
The following theorem allows us to solve the restricted problem of selecting a slowly varying sequence of models (Problem (5)) in polynomial time (proof in Appendix A):
Theorem 3.1.
@@
The restricted formulation (Problem (5)) is an approximation of the full formulation (Problem (1)). To assess the quality of the approximation and the impact of the size of $\mathcal{F}_{b}$ (number of candidate models per batch) on the restricted formulation, we train slowly varying sequences of linear regression models on synthetic data generated according to $y=\boldsymbol{x}^{\top}\boldsymbol{\beta}_{b}+\epsilon$, where, for each batch, the data $X\in\mathbb{R}^{n\times p}(\text{with }n=100,p=10)$ is drawn from $\boldsymbol{x}\sim\mathcal{N}(\boldsymbol{0},\Sigma),\epsilon\sim\mathcal{N}(0,\sigma)$ and the regression coefficients satisfy $|\boldsymbol{\beta}_{b}-\boldsymbol{\beta}_{0}|\leq 0.3$ for some $\boldsymbol{\beta}_{0}\in\{-1,0,1\}$. Note that, thanks to the convexity of the linear regression optimization problem, by varying the value of $\epsilon_{b}$ in Problem (1), we recover the full empirical Pareto frontier. Figure 1 presents the empirical (in-sample) Pareto frontier obtained by solving the full formulation (Problem (1)) as well as the restricted formulation (Problem (5)) for increasing number of candidate models per batch. We also plot the respective true (out-of-sample) Pareto frontiers—visualizing the generalization properties of the proposed approach.
@@
We summarize the key findings from Figure 1:
  • Convexity of the Pareto Frontier: The Pareto frontier is convex, meaning that a small increase in predictive power loss leads to a significant reduction in stability loss. This demonstrates the potential for controlled trade-offs between performance and stability.
    Specifically, to assess the effect of the accuracy tolerance $\alpha$ on the stability gains, we examine the full model’s in-sample curve. We evaluate accuracy tolerances $\alpha\in\{0,0.01,0.03,0.1,0.3\}$, meaning that the predictive power validation loss is constrained to be within $\alpha$ of the best independent model for each batch. The corresponding gains in stability loss range between 2-12%, confirming that moderate increases in predictive loss can substantially enhance stability.
  • Comparison of Full vs. Restricted Models: The restricted model does not always recover PO solutions due to finite model sampling per batch, potentially missing the optimal sequence for certain accuracy tolerances. Consequently, the restricted model finds solutions that are Pareto-dominated by those of the full model. However, increasing $|\mathcal{F}_{b}|$ improves the approximation, allowing the restricted model to better approach the full model’s solutions.
    • For small accuracy tolerances ($\alpha\leq 0.03$), which are the most relevant in practice, the restricted models are 1-3% worse than the full model in both predictive power and stability.
    • For higher accuracy tolerances, the restricted model fails to recover solutions where its stability loss matches that of the full model. This limitation arises because the restricted model explores a finite subset of candidates per batch, making it harder to find sequences with minimal structural change. Achieving stability loss comparable to the full model would require impractically large candidate sets $|\mathcal{F}_{b}|$, effectively ensuring that consecutive batches contain nearly identical models.
  • Generalization Performance: The in-sample vs. out-of-sample predictive power gap is around 20%. Notably, the gap between the full and restricted models remains unchanged between in-sample and out-of-sample settings, validating the theoretical generalization guarantees. Most importantly, the observed difference between in-sample and out-of-sample losses empirically supports the quality of the empirical Pareto frontier as a reasonable approximation of the true Pareto frontier.
@@ 3.2 Adaptive Model Retraining
The intra-sequence stability loss is defined as the sum of pairwise distances of adjacent models of one sequence, i.e., $\mathcal{L}_{S}(\boldsymbol{f})=\sum_{b=1}^{B-1}d(f_{b},f_{b+1})$.
Definition 3.3.
@@
The inter-sequence stability loss is defined as the distance between the final models of two sequences of lengths $B$ and $B^{\prime}$ respectively, i.e., $\mathcal{L}_{S}(\boldsymbol{f},\boldsymbol{f^{\prime}})=d(f_{B},f^{\prime}_{B^{\prime}})$.
@@
For the inter-sequence stability loss, we could alternatively compare all pairs of models in $\boldsymbol{f},\boldsymbol{f^{\prime}}$. However, such an approach (i) does not directly apply to sequences of different lengths and (ii) ignores the fact that we mostly care about the final model from each sequence (which is used in production). The next result promises that, under certain conditions, intra-sequence stability implies inter-sequence stability. In particular, we consider the implications of training a sequence of models that minimizes the intra-sequence stability loss on the inter-sequence stability when a new batch of data is introduced, and the entire sequence is retrained. We work under the additional assumption that the initial model is common in both sequences, i.e., $f_{1}=f_{1}^{\prime}=f_{0}$. This is a realistic assumption as, in practice, $f_{0}$ can be the model that was previously used in production. Further, this assumption can easily be incorporated as an extra constraint in the proposed optimization formulation.
Theorem 3.4.
@@
Let $\boldsymbol{f},\boldsymbol{f^{\prime}}$ be two sequences of models trained using data $D_{1},\dots,D_{B}$ and $D_{1},\dots,D_{B},D_{B+1}$, respectively, and a stability loss metric $d:\mathcal{F}\times\mathcal{F}\mapsto\mathbb{R}_{\geq 0}$. Assume (i) their intra-sequence stability losses satisfy $\mathcal{L}_{S}(\boldsymbol{f})\leq\epsilon$ and $\mathcal{L}_{S}(\boldsymbol{f^{\prime}})\leq\epsilon^{\prime}$, and (ii) the initial models satisfy $f_{1}=f_{1}^{\prime}=f_{0}$. Then, $\mathcal{L}_{S}(\boldsymbol{f},\boldsymbol{f^{\prime}})\leq\epsilon+\epsilon^{\prime}$.
4 Measuring the Stability Loss
To measure the stability loss $\mathcal{L}_{s}(\boldsymbol{f})$ of a sequence $\boldsymbol{f}$, we need to compute the distance $d(f_{i},f_{j})$ between pairs of models $f_{i}$ and $f_{j}$. We consider two primary approaches: structural distances and analytical insight-based distances.
4.1 Structural Distances
This approach measures the distance by directly comparing the models’ structures. Thus, structural distances are model dependent. We discuss structural distances for three model families: regression models, decision tree-based models, and neural networks.
Regression Models. We compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\beta}_{i}-\boldsymbol{\beta}_{j}\|_{2}^{2}$, where $\boldsymbol{\beta}_{i}$ and $\boldsymbol{\beta}_{j}$ are the regressions coefficient vectors of $f_{i}$ and $f_{j}$, respectively. We focus on linear and logistic regression—though our approach directly applies to generalized linear models as well.
Decision Tree-based Models. We use the decision tree distance metric proposed by Bertsimas & Digalakis Jr (2023) that relies on optimally matching the paths in the two trees $\mathbb{T}_{i}$ and $\mathbb{T}_{j}$, denoted as $\mathcal{P}(\mathbb{T}_{i})$ and $\mathcal{P}(\mathbb{T}_{j})$. The distance computation is as follows: first, we calculate the pairwise distance between each path in the first tree and each path in the second tree $d(p,q)$; this pairwise distance takes into account the differences in feature ranges in the two paths as well as the class labels associated with them. Then, we determine the overall distance between two decision trees by finding the optimal matching of these paths, a process that can be computed in polynomial time as a variant of the bipartite matching problem. Formally, assuming without loss of generality that $|\mathcal{P}(\mathbb{T}_{i})|\geq|\mathcal{P}(\mathbb{T}_{j})|$, and given decision variables $x_{pq}=\mathbb{I}$(path $p$ in $\mathbb{T}_{i}$ is matched with path $q$ in $\left.\mathbb{T}_{j}\right)$, the distance is defined as the following (polynomially solvable) linear optimization problem:
[…]
$\quad\sum_{q\in\mathcal{P}(\mathbb{T}_{j})}x_{pq}=1,\forall p\in\mathcal{P}(\mathbb{T}_{i}),\ \sum_{p\in\mathcal{P}(\mathbb{T}_{i})}x_{pq}\geq 1,\forall q\in\mathcal{P}(\mathbb{T}_{j}).$
Neural Networks. To measure the structural difference between neural networks, we compare the singular values of their weight matrices. The intuition behind this approach, inspired by Raghu et al. (2017), is that the singular values of a weight matrix determine how the network transforms input data, describing its ability to stretch, compress, or rotate different feature directions. For each pair of neural networks $f_{i},f_{j}$, we compute the spectral distance by summing the differences in singular values across all corresponding weight matrices: $d(f_{i},f_{j})=\sum_{l}\|\sigma(W_{i}^{(l)})-\sigma(W_{j}^{(l)})\|_{2},$ where $W_{i}^{(l)}$ and $W_{j}^{(l)}$ are the weight matrices of layer $l$ in networks $f_{i}$ and $f_{j}$ and $\sigma(W)$ denotes the ordered singular values of matrix $W$. This metric captures how similarly two networks process data at each layer: small spectral distance implies similar feature transformations, whereas large spectral distance suggests significant differences in how information is propagated.
For comparing neural networks with similar architectures but differing layer sizes, one approach is padding weight matrices with zeros to ensure dimensional consistency before computing singular values. However, when architectures differ significantly, designing a robust distance metric is challenging; we leave the development of such methods for future work.
4.2 Analytical Insight-based Distances
Feature importance is a widely used tool for assessing the analytical insights provided by a model, as it helps explain which features drive predictions. To quantify the stability of these insights across retraining iterations, we measure the distance between models by comparing their feature importance vectors. Given a model $f$, let $\boldsymbol{\phi}_{f}\in\mathbb{R}^{p}$ denote the importance assigned to each of the $p$ features. We compute the distance as $d(f_{i},f_{j})=\|\boldsymbol{\phi}_{f_{i}}-\boldsymbol{\phi}_{f_{j}}\|_{2}^{2}.$
This approach is applicable to both interpretable and black-box models. E.g., for decision tree-based models, we leverage model-specific feature importance metrics such as Gini importance, which measures the contribution of a feature to reducing node impurity, and Gain, which quantifies improvements in classification accuracy from splits made by each feature Breiman et al. (1984); Pedregosa et al. (2011). More generally, for any model type—including black-box models—we can compute feature importance distances using SHAP (SHapley Additive exPlanations) Lundberg & Lee (2017), which offer a standardized methodology to measure feature importance by calculating the average marginal contribution of each feature to the prediction. This provides a unified, model-agnostic way to compare analytical insights across model types.
4.3 Comparing Distance Types
In Appendix C, we analyze the relationship between structural stability, analytical insights, and predictive consistency. We find that the proposed structural distances correlate with both analytical insight-based distances and differences in predictions, confirming that shifts in a model’s learned structure translate into changes in feature attribution and output stability. Logistic regression and decision trees exhibit strong alignment between structural and analytical insight-based stability, whereas neural networks show greater structural variation with lower feature importance shifts, suggesting a more complex relationship between learned weights and interpretability. Across all models, we observe that low structural and analytical insight-based distances often imply similar predictions—but the reverse is not necessarily true; this suggests that the proposed stability loss metrics offer stronger guarantees compared to common prediction-based ones.